You head north from Battle Ground on your motorcycle at 75 kph and two hours later your friend heads south on his motorcycle at 85kph. When the two of you are 710 miles apart, how long has your friend been riding?
Answer by checkley77(12569)

D=RT
710=75T+85(T-2)
710=75T+85T-170
160T=710+170
160T=880
T=880/160
T=5.5 HOURS ANS. FOR THE FIRST MOTORCYCLIST.
5.5-2=3.5 HOURS FOR THE SECOND CYCLLIST.
PROOF:
710=75*5.5+85*3.5
710=412.5+297.5
710=710